A cannonball with a mass of 12.4 kg is fired from a cannon that has a mass of 758.2 kg. The cannonball is fired with a speed of 195.3 m/s, and it takes 0.007 seconds for the cannonball to leave the cannon.



Calculate the net force acting on the cannon.ework-answers/physics/other
1
Expert's answer
2020-01-31T09:56:10-0500
"p_{cannon}=p_{cannonball}=p=mv"

"F=\frac{p}{t}=\frac{12.4\cdot 195.3}{0.007}=345960\ N"